Three years ago Jane was three times as old as Billy. Now she is only twice as old as him. How old is Billy now?
Answer by stanbon(75887)

Three years ago Jane was three times as old as Billy. Now she is only twice as old as him. How old is Billy now?
----
Equations:
j-3 = 3(b-3)
j = 2b
-----
Substitute for "j" and solve for "b":
2b-3 = 3b-9
b = 6 (Billy's age now)
---
Solve for "j":
j = 2b = 12 (Jane's age now)
----------------------------------
Cheers,
Stan H.
=============